# Angle-dependent planar thermal Hall effect by quasi-ballistic phonons in black phosphorus

### v2

@@ Acknowledgements.
Figure S2: Thermal Hall angle in different insulators. The transverse thermal conductivity divided by magnetic field as a function of longitudinal thermal conductivity in different insulators (source: [1, 50, 6, 7, 8, 9, 11, 12, 14, 15, 17, 32]). Note that magnetic insulators with a non-linear field dependence of the thermal Hall effect, such as Lu2V2O7[2] and Cu(1-3, bdc)[3], are not included in the figure.
@@
S5 The universality in thermal Hall physics
Figure S2 illustrates the thermal Hall angle across various insulators. It presents the ratio of transverse thermal conductivity to magnetic field as a function of longitudinal thermal conductivity for these insulators. It is important to note that the figure exclusively features insulators exhibiting a linear or sub-linear relationship between transverse thermal conductivity and magnetic field. Magnetic insulators with a non-linear dependence, such as Lu2V2O7[2] and Cu(1-3, bdc)[3], are not included in the figure.

### v3

@@ VII Author contributions
Figure S2: Phonon spectrum in black phosphorus. Adapted from [34].
Mode Calc. Machida et al. Exp. Fujii et al.
$v_{x}$ LAx 0.454 0.46
TAz 0.490 0.46
TAy 0.119
Average 0.354
$v_{z}$ LAz 0.833 0.96
TAx 0.488
TAy 0.287
Average 0.536
$v_{y}$ LAy 0.505 0.510
TAx 0.160 0.170
TAz 0.289 0.310
Average 0.318 0.330
Table S1: Phonon velocities in black P. Velocities of the three acoustic phonon modes (in units of 104 m/s) along different high symmetry axes
S5 Phonon velocities and phonon spectrum in black Phosphorus.
Table S1 lists the velocity of longitudinal (LA) and transverse acoustic (TA) modes in black P according to the calculated phonon spectrum [34] and the neutron scattering experiments [53]. Note the good agreement between theory and experiment. Remarkably, the longitudinal mode is faster than the transverse mode.
Fig. S2 shows the phonon spectrum up to 500 cm-1. The lowest optical phonon frequency of black phosphorus is around 140 cm-1, much larger than the acoustic phonons below 40 cm-1 which are the focus of this work.